OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (solomonzelman):

3 times Ix-8I +15=18, yes?

OpenStudy (solomonzelman):

3|x − 8| + 15 = 18 3|x − 8| = 3

OpenStudy (solomonzelman):

3 times Ix-8I, yes?

OpenStudy (anonymous):

But how does that get me to the answer ? I'm confused

OpenStudy (solomonzelman):

C

OpenStudy (jdoe0001):

\(\bf 3|x - 8| + 15 = 18\implies 3|x - 8| =18-15\implies 3|x - 8| =3\\ \quad \\ |x - 8|=1\implies \begin{cases} +(x - 8)=1\\ \quad \\ \bf -(x - 8)=1 \end{cases}\)

OpenStudy (anonymous):

9 ?

OpenStudy (solomonzelman):

3(1)+15=18 right?

OpenStudy (anonymous):

right.

OpenStudy (solomonzelman):

Glad I helped you!

OpenStudy (solomonzelman):

(this is my own idea of a sprite profile picture, cool isn't it?)